# unPIC: A Geometric Multiview Prior for Image to 3D Synthesis

### v2

@@ B.1 CROCS Rescaling
Recall that we pre-render multiview NOCS images using a fixed reference frame. This reference frame can be labeled using the source-view camera position $(\theta,\phi,r)=(0,0,1)$ that captures the object in its default creator-intended pose at a default camera distance. At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution. To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section B.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $[[\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$. The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
@@
In the upper row of Figure 8(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $[0,1]^{2}$ CROCS reference square. Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang. This is not ideal—the scale of CROCS values which the model needs to predict depends on the reference frame. Since the reference frame (i.e., source camera position) is not actually supplied to the model, the model can only learn to predict the variable scale by overfitting. We aim to avoid any dependence on the reference frame. To this end, we introduce a CROCS rescaling operation to eliminate the dependence (Figure 8(a), lower row).
Taking a concrete example, Figure 8(b) shows a cube-like object from different angles, along with the pre-rendered NOCS images (second row), and canonicalized CROCS maps without and with rescaling (third and fourth rows). Here are two observations from this figure: 1) The front-facing edge of the cube, visible at $\theta^{\prime}=45$ as yellow in the second row, disappears in the third row because the CROCS values exceed the $[0,1]$ range (and the color map is defined on the same range). If we rescale the values in the third row, we see the edge reappears in the fourth row. 2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to $[-0.2,1.2]$. The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 8(a).
@@
(a) The general case (simplified to 2D). Top row: An object that fits the NOCS reference square exactly at $\theta=0$ is no longer bounded by the CROCS reference square at $\theta^{\prime}\in\{15,30,45\}$. Due to object overhang, some CROCS values lie outside $[0,1]$. The extent of the overhang depends on $\theta^{\prime}$. Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$. This ensures the object is once again bounded tightly by $[0,1]^{2}$.
@@
Figure 9: CROCS when varying the camera elevation angle $\phi$. As in Fig. 3, the large camera denotes the source view; the smaller cameras denote 3 of 7 novel views. The white arrow denotes the normal direction w.r.t. the camera plane (drawn as a white circle connecting the cameras). The colored arrow denotes the normal of the CROCS color cube. Left: at the default angle $\phi=0$, each camera is pointed at a particular face of the RGB color cube, and has a consistent color distribution across examples. Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube. The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $-\phi^{\prime}$ in their respective image planes). The color shifts experienced by the four cameras are clearly inconsistent. Right: If we choose not to tilt the cube through $\phi^{\prime}$, all cameras undergo a small and consistent shift in the distribution of colors toward the up-axis color (blue). If the model can infer $\phi^{\prime}$, it can also predict the shift. We follow the Right approach.
B.2 No Canonicalization for Camera Elevation or Distance
In Section B.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$. Here, we discuss the remaining camera extrinsic parameters. Since all cameras are oriented to face the center of the object, we do not need to consider their rotations. In fact, we do not need to canonicalize for changes in the camera distance $r$. It only serves as a zoom parameter to augment the training data.
@@ Appendix C Additional Results
C.1 Comparisons
Figure 10: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 11: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 12: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 13: Additional qualitative comparison on Objaverse-XL holdouts.
We present further qualitative results comparing unPIC with CAT3D and One-2-3-45 (the top-two competing methods) in Figures 10, 11, 12 and 13.
Besides its multiview inconsistency, One-2-3-45 also has the flaw that it rescales any given object to a canonical size in the image plane. This is visible from the model’s source-pose output (top-left cell in each 2x4 output grid); it is resized compared to the source-pose outputs of other methods. Our approach can cope with arbitrarily sized objects in images despite the fact that it relies on a scale-free 3D shape representation (CROCS).
Note that both CAT3D and our method predict gray values at background pixels. But unPIC produces RGBA outputs rather than RGB—the alpha mask hides the gray background pixels after compositing. We compute masks using posthoc background removal for CAT3D. The remaining methods (One-2-3-45 and OpenLRM) output RGB images with white background pixels.
C.2 Decomposing unPIC’s hierarchical error
In Sec 4.3 we reported metrics from the full hierarchical setup, i.e., on the final output of the prior and decoder stacked together. Since we train the prior and decoder modules separately, we can further analyze the following components of the total error:
  1. 1.
    The error of the prior alone. This describes the geometrical inaccuracy (when the intermediate representation is CROCS) in predicting the 3D shape and pose of a given object.
  2. 2.
    The error of the decoder when it is fed ground-truth CROCS rather than the output of the prior. This describes the difficulty of rendering the object (e.g., predicting the object’s texture) at novel views when extrapolating from a single source image.
We report these components along with total hierarchical error in Table 5. We find that the prior’s error (column 1) is nearly twice that of the oracle decoder (column 2). This highlights the importance of getting the geometry correct. Improving the geometrical prior would help reduce the total error (column 3) more than improving the final rendering (i.e., the decoder module). With geometry predictions as good as the ground-truth CROCS, our full multiview-synthesis error could be 4-5x lower (depending on the dataset).
Table 5: A breakdown of unPIC’s hierarchical error (256x256). We report Mean Square Errors (1e-3) for the prior module, the decoder module when using ground-truth CROCS, and the full error running stacked inference. The MSE is calculated on all 7 novel views.
Image
$\rightarrow$ CROCS
GT CROCS
$\rightarrow$ Pixels
Image
$\rightarrow$ CROCS
$\rightarrow$ Pixels
ObjXL 4.58 2.84 10.52
ABO 3.46 1.65 7.60
GSO 2.69 3.09 7.63
Appendix D More Training and Evaluation Details
Dataset. We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories. We use object type labels predicted with high confidence from [24] to expand the LVIS subset. 2) The KIUI subset  [^2] https://github.com/ashawkey/objaverse_filter comprising 101k additional assets. 3) From the Objaverse-XL alignment subset [6], we used a combination of GLB, OBJ, and FBX assets after filtering out certain (a) terrains and HDRI environment maps, (b) layouts and rooms, and (c) textureless FBX assets that rendered as pink. In total, we used 620k assets (Objaverse and Objaverse-XL) for training. In addition, we held out 50k assets randomly sampled from Objaverse-XL after the filtering step.
Model. We assume an RGB source image, but predict RGBA CROCS and RGBA output images. We train both models with a conditioning dropout probability of 0.05. This opens the possibility of using classifier-free guidance (CFG). For all evaluations in the paper, we use a fixed CFG value of 2.0. The prior model is normally conditioned on the timestep and the source image. We apply CFG to the source image. The decoder model is additionally conditioned on the prior’s output CROCS. We only apply CFG to the source image, as the decoder tracks CROCS pointmaps closely even without CFG.
The architecture for the prior and decoder has just under 150M parameters. It takes about 6 minutes to run the prior and decoder each for 2000 denoising steps, with classifier-free guidance, on an A100. As efficiency wasn’t a core goal for this work, we expect more efficient sampling strategies and better architectures could certainly help reduce the runtime. We use an exponential moving average of the model parameters for evaluation, although this is not essential.

### v3

@@ A.2.1 CROCS Rescaling
Recall that we pre-render multiview NOCS images using a fixed reference frame. This reference frame can be labeled using the source-view camera position $(\theta,\phi,r)=(0,0,1)$ that captures the object in its default creator-intended pose at a default camera distance. At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution. To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section A.2.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$. The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
@@
In the upper row of Figure 6(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $0,1]^{2}$ CROCS reference square. Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang. This is not ideal—the scale of CROCS values which the model needs to predict depends on the reference frame. Since the reference frame (i.e., source camera position) is not actually supplied to the model, the model can only learn to predict the variable scale by overfitting. We aim to avoid any dependence on the reference frame. To this end, we introduce a CROCS rescaling operation to eliminate the dependence (Figure 6(a), lower row).
Taking a concrete example, Figure 6(b) shows a cube-like object from different angles, along with the pre-rendered NOCS images (second row), and canonicalized CROCS maps without and with rescaling (third and fourth rows). Here are two observations from this figure: 1) The front-facing edge of the cube, visible at $\theta^{\prime}=45$ as yellow in the second row, disappears in the third row because the CROCS values exceed the $[0,1]$ range (and the color map is defined on the same range). If we rescale the values in the third row, we see the edge reappears in the fourth row. 2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to [0.2,1.2][-0.2,1.2]. The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 6(a).
@@
(a) The general case (simplified to 2D). Top row: An object that fits the NOCS reference square exactly at $\theta=0$ is no longer bounded by the CROCS reference square at $\theta^{\prime}\in\{15,30,45\}$. Due to object overhang, some CROCS values lie outside $[0,1]$. The extent of the overhang depends on $\theta^{\prime}$. Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$. This ensures the object is once again bounded tightly by $0,1]^{2}$.
@@
Figure 7: CROCS when varying the camera elevation angle $\phi$. As in Fig. 3, the large camera denotes the source view; the smaller cameras denote 3 of 7 novel views. The white arrow denotes the normal direction w.r.t. the camera plane (drawn as a white circle connecting the cameras). The colored arrow denotes the normal of the CROCS color cube. Left: at the default angle $\phi=0$, each camera is pointed at a particular face of the RGB color cube, and has a consistent color distribution across examples. Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube. The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $\phi^{\prime}$ in their respective image planes). The color shifts experienced by the four cameras are clearly inconsistent. Right: If we choose not to tilt the cube through $\phi^{\prime}$, all cameras undergo a small and consistent shift in the distribution of colors toward the up-axis color (blue). If the model can infer $\phi^{\prime}$, it can also predict the shift. We follow the Right approach.
A.2.2 No Canonicalization for Camera Elevation or Distance
In Section A.2.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$. Here, we discuss the remaining camera extrinsic parameters. Since all cameras are oriented to face the center of the object, we do not need to consider their rotations. In fact, we do not need to canonicalize for changes in the camera distance $r$. It only serves as a zoom parameter to augment the training data.
@@ A.2.2 No Canonicalization for Camera Elevation or Distance
Appendix B Training and Evaluation Details
B.1 Dataset
We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories. We use object type labels predicted with high confidence from [17] to expand the LVIS subset. 2) The KIUI subset  [^2] https://github.com/ashawkey/objaverse_filter comprising 101k additional assets. 3) From the Objaverse-XL alignment subset [4], we used a combination of GLB, OBJ, and FBX assets after filtering out certain (a) terrains and HDRI environment maps, (b) layouts and rooms, and (c) textureless FBX assets that rendered as pink. In total, we used 620k assets (Objaverse and Objaverse-XL) for training. In addition, we held out 50k assets randomly sampled from Objaverse-XL after the filtering step.
B.2 Model Hyperparameters
[…]
We train the prior and decoder modules with a conditioning dropout probability of 0.08 on the source image. We do not use any dropout on the CROCS images fed to the decoder to ensure it adheres to the geometry. At test time, we use a fixed classifier-free guidance weight of 2.0 on the source image for both modules. We use an exponential moving average of the model parameters for evaluation, although this is not essential.
The unPIC base model has a total of 1.1B parameters across the prior (470M), decoder (470M), and two VAEs (84M each). It takes about 1min 13s to run all model components, with classifier-free guidance over 1000 denoising steps, on an A100 GPU. Note that we did not attempt to improve the runtime, as that was not the core aim of this work. Techniques such as step distillation, or recent advances in Optimal Flow Matching [22] which facilitate straight trajectories for fast denoising, would be readily applicable to unPIC.
B.3 Choice of K=8 Views
To motivate why $K=8$ views are sufficient to produce a good high-level 3D reconstruction, we take a larger number of ground-truth views ($K=24$), and look at the pairwise similarity between them. This can be computed using CLIP embeddings (like our CLIP-MV metric). See Figure 8 for a heatmap of pairwise similarity scores.
The black crosses that occur in the heatmap show the frequency at which rotated views become dissimilar to each other. We see  6 black crosses on either side of the diagonal of the matrix. This number estimates how many views are sufficient to summarize the full 360-degree rotation. Hence, by generating 7 novel views ($K=8$) for any input image, we can capture 3D structure at sufficient granularity for reconstruction.
Note that this analysis does not take into account top and bottom views of the object.
Figure 8: Pairwise cosine similarities of CLIP embeddings for rotations of a given object, averaged across a subset of ObjaverseXL. The x- and y-axes both show 24 views across a 360-degree rotation.
B.4 U-Net Architecture
The denoising network follows a standard design with a downsampling encoder path, a bottleneck middle path, and an upsampling decoder path with skip connections. The exact architecture is as follows, while the hyperparameters are summarized in Table 6:
  • Input: A noisy latent superimage $X$ and a time-step embedding $t_{emb}$ derived from the diffusion time $t$. Any conditioning superimages (e.g., CROCS latents, $Z^{g}$) are concatenated with $X$ along the channel axis.
[…]
  • Middle Stack: This path consists of three Resnet Blocks with Multi-Head Attention Blocks in between them to process the lowest-resolution feature map.
  • Upsampling Stack: This path mirrors the downsampling path. At each level, the feature map is upsampled, concatenated with the corresponding skip connection from the downsampling path, and passed through three Resnet Blocks. A Multi-Head Attention Block is also applied at specified resolutions.
  • Output: The final feature map is processed through a normalization layer, a Swish activation, and a final $\times3$ convolution to produce the denoised latent superimage of the same shape as the input $X$.
Core Components:
  • Resnet Block: A standard residual block containing two $\times3$ convolutional layers, Group Normalization, and Swish activations. The time-step embedding $t_{emb}$ is projected and used to perform a feature-wise affine transformation (scale and shift) on the feature map after the first normalization layer.
  • Multi-Head Attention Block: A standard multi-head self-attention mechanism applied over the spatial dimensions of the feature map. This allows every location in the feature map to integrate information from all other locations, enabling global communication across the different views tiled in the superimage.
Table 6: U-Net Hyperparameters
Parameter Value
embed_dim 256
channel_mult (1, 2, 2, 3, 3)
num_res_blocks 3
per_head_channels 64
unet_dropout 0.0
norm_type ’Group Norm’
B.5 VAE Fine-Tuning Metrics
We fine-tuned the Stable Diffusion 1.4 VAE, on CROCS and RGB images separately, using a combination of losses: Reconstruction, LPIPS, KL-Divergence, and the GAN loss. See Table 7 showing the improvement in each VAE’s metrics after fine-tuning. The CROCS VAE benefits significantly, and outperforms the RGB VAE in terms of accuracy.
Table 7: VAE metrics on ObjaverseXL before and after fine-tuning on (a) CROCS pointmaps and (b) RGB images.
RGB VAE CROCS VAE
Before Fine-Tuning PSNR 37 38
FID 62 16
KL 28005 20823
After Fine-Tuning PSNR 38 51
FID 16 2
KL 87496 35936
B.6 Assessing Multiview Consistency using MET3R
To assess multiview consistency, we attempted to run MET3R, which is a pose-free way to evaluate the consistency of generated novel views or frames of a generated video. Using ground-truth images, we found that even a 45-degree rotation of an object is too significant for MET3R. The metric produced values (see Fig 9) exceeding the expected value of the metric on random noise (around 0.3). We ran with the standard settings, using the MAST3R backbone, DINO16 features, FeatUp upsampling, and the cosine distance.
Figure 9: MET3R as a metric for multiview consistency produced values in excess of those expected from random noise, even on ground-truth images. The 45-degree rotation in each case is too great to facilitate a dense 3D reconstruction.
B.7 Baselines
For single-image-to-multiview synthesis, we compare unPIC with the following baselines:
  1. 1.
  2. 2.
    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views. It uses the pretrained StableDiffusion v1.5 model as its latent diffusion backbone.
  3. 3.
  4. 4.
    InstantMesh (512x512): a geometry-aware, feed-forward framework that generates a 3D mesh from a single image. We use the Large variant of the model. InstantMesh first employs a multi-view diffusion model, Zero123++, to produce a set of 3D-consistent views from the input. These views are then processed by a transformer-based sparse-view Large Reconstruction Model (LRM) that directly outputs a mesh. The model was trained with direct geometric supervision (depths and normals) on the mesh representation using a differentiable iso-surface extraction module (FlexiCubes). We use the frames rendered from the 3D mesh for our evaluation.
  5. 5.
  6. 6.
@@ Appendix C Additional Results
Task 1: Novel-view synthesis. We provide additional multiview comparisons between unPIC and the baselines in Figure 10. We show five examples at eight target views each.
Task 2: 3D reconstruction. We visualize our point clouds and compare them with two baselines in Figure 15. We also show meshes obtained by applying a surface-reconstruction algorithm (Screened Poisson [18]) on our point clouds. We show 40 examples at one novel view each.
In-the-wild generalization. We apply unPIC to arbitrary real-world images in Figure 23. We also cover failure cases in Figure 24.
Figure 10: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
Figure 11: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
Figure 12: Novel views from baselines versus our method (figure continues on next page). We show an example from our Amazon Berkeley Objects test set.
Figure 13: Novel views from baselines versus our method (figure continues on next page). We show an example from our Digital Twin Catalog test set.
Figure 14: Novel views from baselines versus our method (continued). We show an example from our Digital Twin Catalog test set.
(a) Example 1
(b) Example 2
(c) Example 3
Figure 15: 3D reconstruction results from unPIC vs baseline methods (Part 1 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 4
(b) Example 5
(c) Example 6
Figure 16: 3D reconstruction results from unPIC vs baseline methods (Part 2 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 7
(b) Example 8
(c) Example 9
Figure 17: 3D reconstruction results from unPIC vs baseline methods (Part 3 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 10
(b) Example 11
(c) Example 12
Figure 18: 3D reconstruction results from unPIC vs baseline methods (Part 4 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 13
(b) Example 14
(c) Example 15
Figure 19: 3D reconstruction results from unPIC vs baseline methods (Part 5 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 16
(b) Example 17
(c) Example 18
Figure 20: 3D reconstruction results from unPIC vs baseline methods (Part 6 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 19
(b) Example 20
(c) Example 21
Figure 21: 3D reconstruction results from unPIC vs baseline methods (Part 7 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 22
(b) Example 23
(c) Example 24
Figure 22: 3D reconstruction results from unPIC vs baseline methods (Part 8 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
Figure 23: unPIC on in-the-wild images. Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4).
Figure 24: Failure cases on in-the-wild images. Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4). In the first example, unPIC rotates the person in the image plane, likely due to interpreting the source image as an overhead view. In the second example, unPIC predicts plausible poses, but fails to texture the images correctly, likely due to the out-of-distribution case of multiple humans in the scene.